# InceptionXML: A Lightweight Framework with Synchronized Negative Sampling for Short Text Extreme Classification

### v3

@@ 1. Introduction
Figure 1. SyncXML (Ours) hits the sweet spot in terms of performance on the P@1 metric, training time and model size.
@@
Contributions: These insights motivate us to develop a lightweight architecture and investigate the suitability of convolving over the embedding dimensions of the input in contrast to the words for short-text queries. We thus (i) develop InceptionXML, a lightweight CNN-based encoder, which goes against the traditional paradigm (Kim, 2014; Liu et al., 2017; Dai et al., 2018) of convolving over the words dimension in favor of the embedding dimension, (ii) propose an embedding-enhancement module for learning a word-order agnostic representation of the input, making our approach more robust to lack of structure in short-text queries. Towards scaling the encoder to labels in the order of millions, we (iii) highlight and mitigate the drawbacks of dynamic hard-negative mining by creating a fast and computationally inexpensive model training pipeline SyncXML which synchronizes the label-shortlisting and extreme classification tasks. Finally, we (iv) do a comprehensive comparison of our proposed InceptionXML encoder with pre-trained transformer architectures showing transformers are perhaps an overkill for the short-text XML task.
@@ 2. Related Work
Extreme Classification: The majority of initial works focused on designing one-vs-rest (Babbar and Schölkopf, 2017, 2019; Schultheis and Babbar, 2022), tree-based (Prabhu et al., 2018; Chalkidis et al., 2019; Khandagale et al., 2020) or label embedding based (Bhatia et al., 2015) classifiers with fixed features in the form of bag-of-words representation. With advances in deep learning, jointly learning label and input text spaces has also been developed (Tang et al., 2015; Wang et al., 2018). For XML tasks, recent techniques based on attention mechanism (You et al., 2019) and pre-trained transformer models (Chang et al., 2020; Ye et al., 2020; Jiang et al., 2021; Yu et al., 2020) have shown great promise. For text classification, while (Wang et al., 2017) extended (Kim, 2014) for short input sequences, (Liu et al., 2017) built upon the same for XML tasks.
Short-text Extreme Classification: In XML tasks where the inputs are short text queries, there has been a slew of recent works. Based on the availability of label meta-data, these works can be divided into two categories: (i) ones which make no assumptions regarding label text, i.e., labels are numeric identifiers, such as Astec (Dahiya et al., 2021b) and (ii) others which assume that the labels are endowed with clean label text which include Decaf (Mittal et al., 2021a), GalaXC (Saini et al., 2021), Eclare (Mittal et al., 2021b), and SiameseXML (Dahiya et al., 2021a). Even though the additional label meta-data is useful, it is usually only known for only a small subset of all labels. For instance, on AmazonTitles-3M dataset, label text is available only for 1.3 million labels. Further, the former problem setup, which is the focus of this work, makes no assumption about label-text, and hence is a harder, more general and widely applicable.
@@ 3.2. Embedding Enhancement Module
As mentioned in section 1, the sequence of words in the short-text extreme classification setting do not necessarily observe the syntax of a written language and the presence of a word is more important than its position (Wang and Wang, 2016). Therefore, creating a model agnostic to the input word order is desirable. There exist pooling techniques to make the input word order agnostic, for example, (Dahiya et al., 2021b) use a tf-idf weighted sum of word embeddings as an input to Astec. However, pooling, especially max pooling, often results in loss of information. In a paradigm, where the input to the model is a sequence of 300-dimensional embeddings of 3-5 words, the information is already very scarce and needs alternate methods to be made word order agnostic. To alleviate this, we propose an embedding enhancement module.
@@ 4. SyncXML Training Pipeline
00101010102020202030303030404040400020202020404040406060606080808080100100100100EpochsTraining P@1
@@ 4.1. Synchronized Architecture
4undefundeffor x, y in data:
5undefundefundefundefz = E(x)
6undefundefundefundefh = encoder(z)
7undefundefundefundefy_meta = meta_classifier(h)
8undefundefundefundefy_cluster = label_to_cluster(y)
9undefundefundefundefmeta_loss = bce(y_meta, y_cluster)
@@
11undefundefundefundef# shortlisting top K clusters
12undefundefundefundeftop_k = get_top_K_clusters(y_meta, k)
13undefundefundefundefcandidates = cluster_to_label(top_k)
14undefundefundefundef# add missing positive labels
15undefundefundefundefcandidates = add_missing(candidates,y)
@@
17undefundefundefundef# detached training
18undefundefundefundefif epoch <= epochs/4:
19undefundefundefundefundefundefh = h.detach()
20undefundefundefundefy_ext = ext_classifier(h, candidates)
21undefundefundefundefext_loss = bce(y_ext, y, candidates)
22undefundefundefundefloss = meta_loss + ext_loss
23undefundefundefundefloss.backward()
@@
25undefundefundefundef# gradient descent
26undefundefundefundefupdate(E, encoder, meta_classifier, ext_classifier)
@@ 5.2. Ablation Results
Figure 5. The boxplot shows a variation in performance metrics across 10 runs due to randomly shuffling embedding dimensions before start of training for AmazonTitles-670K dataset. Different scores and statistics can be obtained by adding the values in the y-axis to the base scores on the x-axis.
@@ 6.1. Computational Cost
I. Training time: As shown in Table 8, training time of SyncXML ranges from 4.3 hours on WikiSeeAlsoTitles-350K & AmazonTitles-670K datasets to 27.2 hours on AmazonTitles-3M. We observe a ~$44\%$ decrement in training time by scaling our encoder in the SyncXML training pipeline as compared to the unscaled InceptionXML. As expected, our models train much faster than transformer based approaches (LightXML, APLC-XLNet) while being comparable with Astec.

### v4

@@ 1. Introduction
Figure 1. SyncXML (Ours) hits the sweet spot in terms of scores on the P@1 metric, training time and model size.
@@
Contributions: These insights motivate us to develop a lightweight architecture and investigate the suitability of convolving over the embedding dimensions of the input in contrast to the words for short-text queries. We thus (i) develop InceptionXML, a lightweight CNN-based encoder, which goes against the traditional paradigm (Kim, 2014; Liu et al., 2017; Dai et al., 2018) of convolving over the words dimension in favor of the embedding dimension, (ii) propose an embedding-enhancement module for learning a word-order agnostic representation of the input, making our approach more robust to lack of structure in short-text queries. Towards scaling the encoder to labels in the order of millions, we (iii) highlight and mitigate the drawbacks of dynamic hard-negative mining by creating a fast and computationally inexpensive model training pipeline SyncXML which synchronizes the label-shortlisting and extreme classification tasks. Finally, we (iv) do a comprehensive comparison of our proposed InceptionXML encoder with pre-trained transformer baselines, showing transformers are perhaps an overkill for the short-text XML task.
@@ 2. Related Work
Extreme Classification: The majority of initial works focused on designing one-vs-rest (Babbar and Schölkopf, 2017, 2019; Schultheis and Babbar, 2022), tree-based (Prabhu et al., 2018; Chalkidis et al., 2019; Khandagale et al., 2020) or label embedding based (Bhatia et al., 2015) classifiers with fixed features in the form of bag-of-words representation. With advances in deep learning, jointly learning label and input text spaces has also been developed (Tang et al., 2015; Wang et al., 2018). For XML tasks, recent techniques based on attention mechanism (You et al., 2019) and pre-trained transformer models (Chang et al., 2020; Ye et al., 2020; Jiang et al., 2021; Yu et al., 2020) have shown great promise. For text classification, while (Wang et al., 2017) extended (Kim, 2014) for short input sequences, (Liu et al., 2017) built upon it for XML tasks.
Short-text Extreme Classification: In XML tasks where the inputs are short text queries, there has been a slew of recent works. Based on the availability of label meta-data, these works can be divided into two categories: (i) ones which make no assumptions regarding label text, i.e., labels are numeric identifiers, such as Astec (Dahiya et al., 2021b) and (ii) others which assume that the labels are endowed with clean label text which include Decaf (Mittal et al., 2021a), GalaXC (Saini et al., 2021), Eclare (Mittal et al., 2021b), and SiameseXML (Dahiya et al., 2021a). Even though the additional label meta-data is useful, it is usually only known for only a small subset of all labels. For instance, on AmazonTitles-3M dataset, label text is available only for 1.3 million labels. Further, the former problem setup, which is the focus of this work, makes no assumption about label-text, and hence is harder, more general and widely applicable.
@@ 3.2. Embedding Enhancement Module
As mentioned in section 1, the sequence of words in the short-text extreme classification setting do not necessarily observe the syntax of written language, and the presence of a word is more important than its position (Wang and Wang, 2016). Therefore, creating a model agnostic to the input word order is desirable. There exist pooling techniques to make the input word order agnostic, for example, (Dahiya et al., 2021b) use a tf-idf weighted sum of word embeddings as an input to Astec. However, pooling, especially max pooling, often results in loss of information. In a paradigm, where the input is a sequence of 300-dimensional embeddings of 3-5 words, the information is already very scarce and needs alternate methods to be made word order agnostic. To alleviate this, we propose an embedding enhancement module.
@@ 4. SyncXML Training Pipeline
00101010102020202030303030404040400020202020404040406060606080808080100100100100EpochsTraining P@1LightXML extremeLightXML metaSyncXML extremeSyncXML meta
@@ 4.1. Synchronized Architecture
4 for x, y in data:
5 z = E(x)
6 h = encoder(z)
7 y_meta = meta_classifier(h)
8 y_cluster = label_to_cluster(y)
9 meta_loss = bce(y_meta, y_cluster)
@@
11 # shortlisting top K clusters
12 top_k = get_top_K_clusters(y_meta, k)
13 candidates = cluster_to_label(top_k)
14 # add missing positive labels
15 candidates = add_missing(candidates,y)
@@
17 # detached training
18 if epoch <= epochs/4:
19 h = h.detach()
20 y_ext = ext_classifier(h, candidates)
21 ext_loss = bce(y_ext, y, candidates)
22 loss = meta_loss + ext_loss
23 loss.backward()
@@
25 # gradient descent
26 update(E, encoder, meta_classifier, ext_classifier)
@@ 5.2. Ablation Results
Figure 5. The boxplot shows a variation in performance metrics across 10 runs by randomly shuffling embedding dimensions before start of training for AmazonTitles-670K dataset. Different scores and statistics can be obtained by adding the values in the y-axis to the base scores on the x-axis.
@@ 6.1. Computational Cost
I. Training time: As shown in Table 8, training time of SyncXML ranges from 4.3 hours on WikiSeeAlsoTitles-350K & AmazonTitles-670K datasets to 27.2 hours on AmazonTitles-3M. We observe a ~$44\%$ decrement in training time by scaling our encoder in the SyncXML training pipeline as compared to the unscaled InceptionXML. As expected, our models train much faster than transformer baselines (LightXML, APLC-XLNet) while being comparable with Astec.